The slope of a line is 3/4 and its y-intercept is (0,6). What is its x-intercept?
Answer by oscargut(2103)

Equation is: (3/4)x+6
to find the x-intercept we must solve:
(3/4)x+6 = 0
(3/4)x = -6
3x = -24
x = -8
Answer: x-intercept is -8 (or (-8,0))

I can help you with more questions at: mthman@gmail.com
Thanks